The cost, C, of manufacturing x water sprinklers is given by the formula C=x^2-27x-20. Determine the number of water sprinklers manufactured at a cost of $70.
Answer by edjones(8007)

x^2-27x-20=70
x^2-27x-90=0
(x-30)(x+3)=0
x=30 sprinklers.
.
Ed